OpenStudy (maddy1251):

Need help with this problem?? A 10.0 kg super ball falling at 10.0 m/s hits a hard surface and bounces off the floor going 8 m/s. The ground exerts a force on the ball of 720 N. How long was the ball in contact with the ground? 0.75 s 1.0 s 0.25 s 0.50 s I guessed 0.25 s.

OpenStudy (nino_jim):

whats the formula

OpenStudy (maddy1251):

@Nino_jim I am trying to find one.

OpenStudy (nino_jim):

@Maddy1251 well tell me when you get it

OpenStudy (maddy1251):

@Michele_Laino Do you have an idea?

OpenStudy (maddy1251):

@Nino_jim the answer is 0.25 s We take initial v and final v (10 m/s x 10 = 100 but going downward in v is -100 and fv is 8 x 10 = 80 and is upward so +80 --> +80 - (-180) = 180 and then divide by each answer choice to get 720 N which is (180/0.25)= 720 N.